OpenStudy (anonymous):

Please help. ----------- Problem below! The height is multiplied by 3. Then the area increases by 3*3=9 So would it be 3x or are they asking for the final product which is 9? @ganeshie8 @amistre64

OpenStudy (anonymous):

@Nurali That's where I'm confused- are they asking about when it's "tripled".. or the final product "nine"? I'm confused! @ganeshie8 @amistre64

OpenStudy (anonymous):

@ganeshie8 and @amistre64 I am confused because they are both options and I'm not positive what part they are asking for.

OpenStudy (amistre64):

surface area: 4(6n*9n)/2 + (9n)^2 let n=1, and let n=3 and compare the results 4(54)/2 + 81 = 189 4(18*27)/2 + (27)^2 = 1701 so its going to be bigger by 1701/189

OpenStudy (anonymous):

= 9

OpenStudy (amistre64):

logically, the area will be bigger; so 3 or 9 times bigger; and since area is affected more than linear, 3times on the linear will be more than 3times on the area

OpenStudy (amistre64):

but yes, mathically it comes to 9

OpenStudy (anonymous):

In other words , you are saying it is either 3 or 9x bigger; given to the calculations we've done, it can only be nine times the original pyramid’s surface area, correct?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

Thanks, I get it now :)

OpenStudy (amistre64):

good luck :)